OpenStudy (anonymous):

if y=3x+2cos x, find dy/dx and hence show that y increases as x increases. i know how to get dy/dx but i dont know how to prove that 'y increases as x increases' please help!

OpenStudy (zehanz):

You will have found then that dy/dx=3-2sinx. Take a good look at this formula. What are the possible outcomes (the range) of dy/dx?

OpenStudy (anonymous):

it is pi/6 or - pi/6 i think?!?!?

OpenStudy (zehanz):

No, you should remember that you can put any number x into dy/dx=3-2sinx. If you do this, what are all the possible outcomes of 3-2sinx? Maybe first: what are the possible outcomes of sinx?

OpenStudy (anonymous):

lowest x value to higher one?

OpenStudy (zehanz):

You know the range of sinx: it is [-1,1]. Whatever x to put into the sine function, the outcome is always a number between -1 and 1 (included). So if you calculate 2sinx, the outcome of that is always a number in the interval [-2,2]. Are you with me so far?

OpenStudy (anonymous):

ahhh true..

OpenStudy (anonymous):

i forgot about the restrictions

OpenStudy (zehanz):

OK, now we have this number in the interval [-2,2]. To calculate dy/dx, you subtract the number from 3, after all, dy/dx=3-2sinx. To see what the possible outcomes of dy/dx are, look at the extremes -2 and 2: 3--2=5 3-2=1 So...dy/dx alis always in the interval [1,5]. That means: dy/dx is always positive.

OpenStudy (anonymous):

so.. 2sinx could be replaced with -2 or 2?

OpenStudy (zehanz):

Also, you know that dy/dx is the slope of the tangent line to the graph of y=3x+2cosx. So the slope is always a positive number, so y is always increasing. Remember: y' positive means y increasing y' negative means y decreasing

OpenStudy (zehanz):

No you cannot replace 2sinx by 2 or -2, it is just that I want to see what values 3-2sinx can take. At the very LEAST it gets to 3-2 = 1 At the MOST it gets to 3--2 = 5, so the range is [1,5]

OpenStudy (anonymous):

Your statement is very easy! thank you so much for the help!

OpenStudy (zehanz):

Happy New Year!

OpenStudy (anonymous):

you too!! :)